OpenStudy (anonymous):

express 9-y^2 as the product of two binomial factors

OpenStudy (ajprincess):

Difference of squares. 9=3^2 a^2-b^2=(a-b)(a+b)

OpenStudy (ajprincess):

here a=3 and b=y

OpenStudy (anonymous):

(3-y)(3+y)?

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

welcome:)